OpenStudy (katherinesmith):

The time required to finish a test in normally distributed with a mean of 80 minutes and a standard deviation of 15 minutes. What is the probability that a student chosen at random will finish the test in more than 95 minutes?

OpenStudy (blockcolder):

Basically, you want \(P(X>95)=P(X-80>15)=P\left(\frac{X-80}{15}>1\right)=P(Z>1).\) Consult the back of your stat book so that you can find P(Z>1).